OpenStudy (harsha19111999):

Java Programming. If x=0 and x=x+1, what is the output?

OpenStudy (anonymous):

If `x=0` then `x=x+1` becomes `x=0+1` becomes `x=1`

OpenStudy (harsha19111999):

But what is the output for continuous program? Like if I want to calculate the number of members in a room

OpenStudy (harsha19111999):

@wio

OpenStudy (harsha19111999):

@dan815

OpenStudy (harsha19111999):

@ganeshie8

OpenStudy (harsha19111999):

@Koikkara

OpenStudy (koikkara):

@Harsha19111999 @wio Gave You the hint needed to find your answer...!!

OpenStudy (davidusa):

here is the file. compile it and run. remember, everything is case sensitive. English not english Morse not morse

OpenStudy (harsha19111999):

Can you compile it successfully?

OpenStudy (davidusa):

yes.

OpenStudy (harsha19111999):

Because it is showing me 3 errors

OpenStudy (davidusa):

i know why! u need the input.class file that i have here download and put in same folder as the other file

OpenStudy (harsha19111999):

No I can't compile it.

OpenStudy (davidusa):

@Harsha19111999 hm... it works for me though

OpenStudy (davidusa):

@Harsha19111999 heres the class file of morse code by the way morsecodetest2 is same as morsecode so its not because of that

OpenStudy (davidusa):

@Harsha19111999 look go here http://screencast-o-matic.com/watch/c21qbpnCvU its me!!!!

OpenStudy (davidusa):

@Harsha19111999 come on where are u can u help me look at my video

OpenStudy (harsha19111999):

Yes I saw it. I will get back to you tomorrow. Don't worry

OpenStudy (harsha19111999):

Good night @DavidUsa

OpenStudy (harsha19111999):

I have to sleep. Its 11:30 PM here

OpenStudy (davidusa):

ok bye. @WIO PLEASE HELP ME Yay ok harsha good night

OpenStudy (harsha19111999):

:)

OpenStudy (e.mccormick):

Do not mix questions together like this. It is considered spamming another question. For the original question: Java Programming. If x=0 and x=x+1, what is the output? It is going to depend on how it is coded. Because you used the word if, which has a very specific meaning in code, then that psudocode could result in whatever the FALSE result of the if statement is.